Avarage age at your parents is 45 years. When your age is counted the avarage 33. What is your age?
Answer by stanbon(75887)

Average age at your parents is 45 years.
---
(m+d)/2 = 45
m+d = 90
---------------------
When your age is counted the avarage 33. What is your age?
(m+d+y)/3 = 33
m+d+y = 99
------
Substitute for "m+d" and solve for "y":
90 + y = 99
y = 9 yrs (your age now)
------------------------------
Cheers,
Stan H.
=========================